# Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation Short Title: Electron-irradiated radicals for DNP

### v2

@@ 6.8 Alanine SAR EPR Fitting and EPR Simulation Details
Fitting was performed using a two-stage optimisation approach. First, we employed Matlab’s surrogateopt function with 10,000 maximum function evaluations, exploring a bounded parameter space centred around initial estimates. The objective function combined both spectral and integral matching metrics to ensure proper reproduction of both line shapes and relative intensities. Specifically, we minimised:
@@
where $\mathcal{I}$ represents cumulative integrals of spectra and $\mathcal{S}$ represents the normalised spectra themselves.
The final fit was further refined using Bayesian optimization (bayesopt) with additional iterations (150) to ensure convergence to the global minimum. The fitting achieved satisfactory spectral reproduction with the dominant R1 species, though slight deviations suggest potential minor contributions from R2 and R3 species or other structural heterogeneities in the 13C-labelled samples. Full experimental metadata, simulation code, and raw experimental data are provided as supplemental data S2.
We should note that the alanine radical systems formed following irradiation by ionising radiation have been studied in depth by EPR, and many reported parameters have been obtained for both the initial high-energy primary radical, and three subsequent radical species. A reasonable fit to our data can be obtained by previously reported parameters, (?) shown in Fig. S6, and these are enclosed (with EPR relaxometry data) as Data S2. These fits do not include any degree of crystallographic ordering, as rotating the tube did not affect signal intensity. At low temperature, it is likely that radical species interconvert and in particular may lose direct molecular character, with some degree of between-molecule interactions possible within the molecular sites that are not equivalent in the unit cell. The low temperature data can be modelled to a moderate level using parameters derived from those published for the primary radical, (?) summed with smaller contributions from the known SAR species R1, R2, R3. (?, ?) See code in Data S2 for specific values.
With regards to partial ordering which is invoked in the analysis of the XRD and DNP datasets but not explicitly observed in the EPR data, we should further stress the degree to which sample preparation may affect this parameter. Both DNP and XRD samples are mildly compressed into their containers, either manually when entering the sample chamber (XRD) or by 3 bar helium gas as part of an atmospheric purge cycle (DNP). In contrast, EPR tubes are narrow diameter and it was not readily possible to pack powder samples densely within them, which again might mean that EPR undertaken in these spectrometers is not representative of that within the DNP cryostat. Furthermore, reflecting the presence of different interaction strengths at different field and temperature combinations, the exact microscopic picture may well depend upon parameters such as the rate at which the sample was lowered and frozen in the magnetic field in the DNP polariser. Furthermore, the ordering we report in both the crystal and XRD instrument frames may well be aligned differently to that in our DNP cryostat, itself with $B$ aligned perpendicular to the laboratory floor (unlike EPR spectrometers, where it is parallel). Although the crystallite size distribution from XRD data is consistent with microcrystallites having a narrow size distribution of mean length $580.8\pm 37.6$ nm, and therefore the effect of gravity is expected to be minimal, such considerations cannot be ruled out as influencing the distribution of environments contained within the sample.
Figure S6: A demonstration of simulated literature values for reported alanine radicals. At room temperature, a moderately good fit to (A) natural abundance data can be obtained from a 70%/30% mixture of the radicals SAR1 and SAR2. A similar fit can be obtained for (B) the 13-C labelled alanine powder. At low temperature (C), a freely fitted mixture of R1, R2 and the primary radical can be found to approximate the shape of the experimental data. These data have been field corrected and frequency corrected to 9.6 GHz
@@ 6.9 X-ray diffraction texture analysis
The pole figures were generated for the principal crystallographic directions (normal direction [ND], rolling direction [RD], and transverse direction [TD]) to visualise the orientation distribution. The texture analysis revealed that alanine crystals in glycerol adopt specific preferential orientations, likely due to interactions between the crystal faces and the glycerol matrix during sample preparation. The refined texture model was incorporated into the final Rietveld refinement, which significantly improved the fit to the experimental data, confirming that the observed intensity variations were indeed due to preferred orientation rather than structural factors. These data are provided as supplemental data S1.
@@ 6.10 Molecular Dynamics Simulations
  • Energy minimization: The system was energy-minimised to eliminate bad contacts using LAMMPS.
@@
Post-processing was performed in VMD, (?) and density and the radial distribution function computed at the end of the equilibration process (c.f. Fig. S9). It was found that liquid glycerol did not penetrate substantially into the alanine domain on the timescale of the simulation, suggesting limited miscibility between the two substances under these conditions. The overall radial distribution function $g(r)$ exhibited features consistent with liquid-phase behaviour, including a prominent first coordination shell peak at approximately 2 Å followed by diminishing oscillations that approached uniformity at larger distances. This $g(r)$ profile confirms the absence of long-range crystalline order and indicates a liquid-like molecular arrangement with short-range spatial correlations characteristic of hydrogen-bonded systems. While our simulation approach cannot definitively determine the global thermodynamic minimum of the system (which might involve a crystalline alanine phase separate from glycerol), it effectively establishes the relative stability of the mixed state and characterises the molecular interactions at the interface between alanine and glycerol domains. The persistence of distinct domains throughout the 30 ns simulation provides valuable information about the limited mutual solubility of these compounds under the studied conditions.
@@ 6.11 Spin-dynamics simulations
In an attempt to semi-empirically predict the asymmetric frequency sweep curve, a computational study on a model of the radical-damaged singular alanine crystal was undertaken. The spatial distribution of stable alanine radicals (SARs) was modelled using the known crystallographic data from the L-alanine structure provided by neutron diffraction, assuming these are unchanged by irradiation. Since electron irradiation creates radicals at random positions within the crystal lattice, we employed bootstrap sampling (10,000 iterations) of possible radical positions to determine the statistical distribution of inter-radical distances, written using ASE in Python. (?)
@@
modulation rate $=100~\text{Hz}$ (S10)
@@
The powder averaging was performed using the icos-2ang-12pts grid for the basic simulations and rep-2ang-100pts-sph for more detailed models. The polarisation transfer was monitored using detection operators corresponding to the 13C $L_{z}$ Liouvillian terms. For extended simulations with multiple 13C nuclei, we incorporated 56 carbon positions from a supercell of the alanine unit cell (which contains four molecules), with the radical species positioned at sites 21 and 5 in the lattice. The electron-electron dipolar coupling was calculated from their spatial separation within the crystal structure. Periodic boundary conditions were included. The simulation ran on 200 nodes, each containing 48 core Cascade Lake processors (Intel Xeon Platinum 8268 CPU @ 2.90GHz) with 392 GB of ram, and took approximately 4 days to complete.
@@ 6.12 Analytic models
Thermal mixing is a key mechanism in Dynamic Nuclear Polarisation (DNP), widely taken to be dominant in samples used in clinical practice, and theoretically challenging. To quantitatively assess the DNP polarisation behaviour, we implemented a numerical simulation based on Wenckebach’s framework in the low temperature limit  (?).
In this model, the electron spin system is characterised by two parameters: the inverse electron Zeeman temperature $\alpha=\hbar/k_{B}T_{Z}$ and the inverse electron non-Zeeman temperature $\beta_{NZ}=\hbar/k_{B}T_{NZ}$. This assumption splits the electronic states into those that are polarised thermally (Zeeman) and those that will ultimately transfer energy to the nuclei in the irradiation phase of DNP (non-Zeeman). This results in an electron spin polarisation as a function of resonance frequency:
@@
where $g(\omega)$ is the normalised ESR spectral density with centre of mass $\omega_{0}$, $\omega_{m}$ is the applied microwave frequency, $W(\omega_{m})=\frac{1}{2}\pi\omega_{1S}^{2}g(\omega_{m})$ represents the microwave power, $\omega_{1S}=\gamma_{S}|B_{1}|$ is the electron Rabi frequency, $T_{1S}$ is the electron spin-lattice relaxation time, and $P_{L}=\tanh(\frac{1}{2}\omega_{0}\beta_{L})$ is the thermal equilibrium polarisation with $\beta_{L}=\hbar/k_{B}T_{L}$ being the inverse lattice temperature. High and low temperature limits are defined by the $\omega_{0}\beta\approx 1$ boundary at roughly around 1 K.
@@
These equations were evaluated using adaptive Gaussian quadrature (NIntegrate) with specified domain constraints. To ensure numerical stability, the working precision was set to 20 digits, with recursion limits extended to handle the complex structure of the integrands. This process took approximately one day on a dual Xeon gold workstation with 768 GB ram, and the simulation results are shown in Fig. S11A, with experimental integrated EPR lineshapes shown in Fig. S11B. They produce a bimodal curve qualitatively similar to that in Spinach, and do not reflect experimental results.
@@
As a semi-empirical test with the observed (and not modellable) EPR lineshape obtained, a further simulation was conducted using a semi-empirically scaled EPR lineshape based on that obtained of the alanine/glycerol mixture at 5 K. This simulation was performed in Matlab (providing an independent reimplementation of the model above) and scaled the centre linearly and width (by approximately a factor of two) of the EPR line to be that appropriate at 6.7 T under the assumption that these qualitative features would describe the system, as it is not possible for us to directly measure the EPR spectrum under DNP conditions of 1.4 K and 6.7 T. Both the high temperature and low temperature approximations as defined by Wenckebach have been solved separately and compared to experimental data. Even at 30 kGy irradiation, where a small amount of bimodal behaviour is observed, this model (which accurately reproduces the behaviour of e.g. trityl radicals, TEMPO, Totapol, and other samples) does not fit the data obtained empirically, shown in Fig. S12. We are confident that the bandwidth of the (commercially supplied) microwave excitation source exceeds that shown, and is not the limiting factor in these experiments.
@@ 6.13 QuantumESPRESSO
First-principles electronic structure calculations were performed using the plane-wave pseudopotential approach as implemented in the Quantum ESPRESSO (QE) package (version 7.2)  (?, ?). The Perdew-Burke-Ernzerhof (PBE) generalised gradient approximation  (?) was used for the exchange-correlation functional. Projector augmented-wave (PAW) pseudopotentials were employed for carbon and oxygen atoms, while ultrasoft pseudopotentials were used for hydrogen and nitrogen atoms, all obtained from the Standard Solid-State Pseudopotential (SSSP) library (efficiency version)  (?).
@@
6.14 Products of dissolution
There were no 13C-labelled impurity peaks present in the dissolution products, and the produced hyperpolarised spectra were largely indistinguishable from that using trityl radicals. This is shown in detail below, in Fig. S14.
Figure S14: Hyperpolarised alanine produced either with trityl radical or with e-beam irradiation, both when initially dissolved (1 average, WALTS-16 proton decoupling) and corresponding thermal-equilibrium acquisition (256 averages, fully relaxed, WALTZ-16 proton decoupling). No impurity peaks were observed in either spectra. Data obtained at 1.4 T on a Magritek Spinsolve Ultra. The enhancement factor is approximately 142,000.
7 Data captions
Caption for Data S1.
Caption for Data S2.
Raw EPR data of irradiated and natural abundance alanine used for modelling in this work, together with MATLAB files that reproduce either the modelled spectrum or $T_{1e}$ or $T_{m}$ fits.

### v3

@@ 6.5 Alanine SAR EPR Fitting
Fitting was performed using a two-stage optimization approach. First, we employed MATLAB’s surrogateopt function with 10,000 maximum function evaluations, exploring a bounded parameter space centred around initial estimates. The objective function combined both spectral and integral matching metrics to ensure proper reproduction of both line shapes and relative intensities. Specifically, we minimized:
@@
where $\mathcal{I}$ represents cumulative integrals of spectra and $\mathcal{S}$ represents the normalized spectra themselves.
The final fit was further refined using Bayesian optimization with additional iterations (150) to ensure convergence to the global minimum. The fitting achieved satisfactory spectral reproduction with the dominant R1 species, though slight deviations suggest potential minor contributions from R2 and R3 species or other structural heterogeneities in the 13C-labelled samples.
6.5.1 Code Listing 1
%See doi: 10.1021/jp972158k for details
published_eigenvals=[2.0041 2.0034 2.0024];
published_eigenvecs=[0.828 0.030 0.559;
    -0.428 0.678 0.598;
    0.362 0.735 -0.574];
g_tensor=published_eigenvecs*diag(published_eigenvals)*inv(published_eigenvecs);
%Anisotropic couplings of the alanine R1 radical
A1 = [-87.8 -52.5 -27.8];          % H-alpha principal hfc values (MHz)
A2 = [74.7 67.6 67.3];             % H-beta principal hfc values (MHz)
A3 = [5.1 -2.1 -2.6];              % H-gamma principal hfc values (MHz)
%(The system is [A1;A2;A2;A2;A3]; for the five different protons)
%Euler angles are
V1 =[-0.6420   1.3484    0.5622];  % H-alpha Euler angles
V2 =[0.0574    2.2199    1.9514];  % H-beta Euler angles
V3 =[2.9103    2.5350   -1.6269];  % H-gamma Euler angles
Sys = struct(’S’,1/2,’g’,published_eigenvals);
Sys = nucspinadd(Sys, ’1H’, A1,V1); %HAlpha
Sys = nucspinadd(Sys, ’1H’, A2,V2); %HBeta
Sys = nucspinadd(Sys, ’1H’, A2,V2); %HBeta
Sys = nucspinadd(Sys, ’1H’, A2,V2); %HBeta
Sys = nucspinadd(Sys, ’1H’, A3,V3); %HGamma
%We now add the 13C below and numerically solve for its A and V.
@@ 6.6 X-ray diffraction texture analysis
The pole figures were generated for the principal crystallographic directions (normal direction [ND], rolling direction [RD], and transverse direction [TD]) to visualize the orientation distribution. The texture analysis revealed that alanine crystals in glycerol adopt specific preferential orientations, likely due to interactions between the crystal faces and the glycerol matrix during sample preparation. The refined texture model was incorporated into the final Rietveld refinement, which significantly improved the fit to the experimental data, confirming that the observed intensity variations were indeed due to preferred orientation rather than structural factors. These data are provided as supplemental data S1.
Caption for Data S1.
@@ 6.7 Molecular Dynamics Simulations
  • Energy minimization: The system was energy-minimized to eliminate bad contacts using LAMMPS.
@@
Post-processing was performed in VMD, (?) and density and the radial distribution function computed at the end of the equilibration process (c.f. Fig. S5). It was found that liquid glycerol did not penetrate substantially into the alanine domain on the timescale of the simulation, suggesting limited miscibility between the two substances under these conditions. The overall radial distribution function $g(r)$ exhibited features consistent with liquid-phase behaviour, including a prominent first coordination shell peak at approximately 2 Å followed by diminishing oscillations that approached uniformity at larger distances. This $g(r)$ profile confirms the absence of long-range crystalline order and indicates a liquid-like molecular arrangement with short-range spatial correlations characteristic of hydrogen-bonded systems. While our simulation approach cannot definitively determine the global thermodynamic minimum of the system (which might involve a crystalline alanine phase separate from glycerol), it effectively establishes the relative stability of the mixed state and characterizes the molecular interactions at the interface between alanine and glycerol domains. The persistence of distinct domains throughout the 30 ns simulation provides valuable information about the limited mutual solubility of these compounds under the studied conditions.
@@ 6.8 Spin-dynamics simulations
In an attempt to semi-empirically predict the asymmetric frequency sweep curve, a computational study on a model of the radical-damaged singular alanine crystal was undertaken. The spatial distribution of stable alanine radicals (SARs) was modelled using the known crystallographic data from the L-alanine structure provided by neutron diffraction. Since electron irradiation creates radicals at random positions within the crystal lattice, we employed bootstrap sampling (10,000 iterations) of possible radical positions to determine the statistical distribution of inter-radical distances, written using ASE in Python. (?)
@@
modulation rate $=1~\text{kHz}$ (S6)
@@
The powder averaging was performed using the icos-2ang-12pts grid for the basic simulations and rep-2ang-100pts-sph for more detailed models. The polarization transfer was monitored using detection operators corresponding to the 13C $L_{z}$ Liouvillian terms. For extended simulations with multiple 13C nuclei, we incorporated 56 carbon positions from a supercell of the alanine unit cell (which contains four molecules), with the radical species positioned at sites 21 and 5 in the lattice. The electron-electron dipolar coupling was calculated from their spatial separation within the crystal structure. Periodic boundary conditions were included. The simulation ran on 200 nodes, each containing 48 core Cascade Lake processors (Intel Xeon Platinum 8268 CPU @ 2.90GHz) with 392 GB of ram, and took approximately 4 days to complete.
@@ 6.9 Analytic models
Thermal mixing is a key mechanism in Dynamic Nuclear Polarization (DNP), widely taken to be dominant in samples used in clinical practice, and theoretically challenging. To quantitatively assess the DNP polarization behaviour, we implemented a numerical simulation based on Wenckebach’s framework in the low temperature limit  (?).
In this model, the electron spin system is characterized by two parameters: the inverse electron Zeeman temperature $\alpha=\hbar/k_{B}T_{Z}$ and the inverse electron non-Zeeman temperature $\beta_{NZ}=\hbar/k_{B}T_{NZ}$. This assumption splits the electronic states into those that are polarised thermally (Zeeman) and those that will ultimately transfer energy to the nuclei in the irradiation phase of DNP (non-Zeeman). This results in an electron spin polarization as a function of resonance frequency:
@@
where $g(\omega)$ is the normalised ESR spectral density with centre of mass $\omega_{0}$, $\omega_{m}$ is the applied microwave frequency, $W(\omega_{m})=\frac{1}{2}\pi\omega_{1S}^{2}g(\omega_{m})$ represents the microwave power, $\omega_{1S}=\gamma_{S}|B_{1}|$ is the electron Rabi frequency, $T_{1S}$ is the electron spin-lattice relaxation time, and $P_{L}=\tanh(\frac{1}{2}\omega_{0}\beta_{L})$ is the thermal equilibrium polarization with $\beta_{L}=\hbar/k_{B}T_{L}$ being the inverse lattice temperature. High and low temperature limits are defined by the $\omega_{0}\beta\approx 1$ boundary at roughly around 1 K.
@@
These equations were evaluated using adaptive Gaussian quadrature (NIntegrate) with specified domain constraints. To ensure numerical stability, the working precision was set to 20 digits, with recursion limits extended to handle the complex structure of the integrands. This process took approximately 1 day on a dual Xeon gold workstation with 768 GB ram, and the simulation results are shown in Fig. S7A, with experimental integrated EPR lineshapes shown in Fig. S7B. They produce a bimodal curve qualitatively similar to that in Spinach, and do not reflect experimental results.
@@
As a semi-empirical test with the observed (and not modellable) EPR lineshape obtained, a further simulation was conducted using a semi-empirically scaled EPR lineshape based on that obtained of the alanine/glycerol mixture at 5 K. This simulation was performed in Matlab (providing a novel reimplementation of the model above) and scaled the centre linearly and width (by approximately a factor of two) of the EPR line to be that appropriate at 6.7 T under the assumption that these qualitative features would describe the system, as it is not possible for us to directly measure the EPR spectrum under DNP conditions of 1.4 K and 6.7 K. Both the high temperature and low temperature approximations as defined by Wenckebach have been solved separately and compared to experimental data. Even at 30 kGy irradiation, where a small amount of bimodal behaviour is observed, this model (which accurately reproduces the behaviour of e.g. trityl radicals, TEMPO, Totapol, and other samples) does not fit the data obtained empirically, shown in Fig. S8. We are confident that the bandwidth of the (commercially supplied) microwave excitation source exceeds that shown, and is not the limiting factor in these experiments.
@@ 6.10 QuantumESPRESSO
First-principles electronic structure calculations were performed using the plane-wave pseudopotential approach as implemented in the Quantum ESPRESSO (QE) package (version 7.2)  (?, ?). The Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation  (?) was used for the exchange-correlation functional. Projector augmented-wave (PAW) pseudopotentials were employed for carbon and oxygen atoms, while ultrasoft pseudopotentials were used for hydrogen and nitrogen atoms, all obtained from the Standard Solid-State Pseudopotential (SSSP) library (efficiency version)  (?).